### v1

@@ 4.1 Data
For the study region, we focus on the South American continent and the surrounding oceans. Specifically, the targeted area spans from latitude 0°N to 63°S and from longitude -90°W to -27°E. The training period comprises ERA5 data from 1992-01-01 to 2011-01-01. The range of years included for evaluation on ERA5 and GFDL spans from 2011-01-02 to 2014-12-01.
@@
The climate model output is taken from a state-of-the-art ESM from Phase 6 of the Coupled Model Intercomparison Project (CMIP6), namely the GFDL-ESM4 [3]. The dataset contains daily precipitation data of the first ensemble member (r1i1p1f1) of the historical simulation (esm-hist). The data is available from 1850 to 2014, at 1° latitudinal and 1.25° longitudinal resolution and daily temporal resolution.
@@
In order to benchmark our method, we first apply bi-linear interpolation to increase the resolution of the GFDL fields from 1° to 0.25°. After that, we apply Quantile delta mapping [1] to fit the upsampled GFDL data to the original 0.25° ERA5 data. QM is fitted on past observations, and can then be used to correct the statistics of any (past/present) ESM field towards that reference period. We use quantile delta mapping and chose the training period of ERA5 from 1992-01-01 to 2011-01-01 as the reference period for fitting GFDL to ERA5. The benchmark dataset to evaluate our approach is then constructed by applying QM to the validation period of GFDL (2011-01-02 to 2014-12-31).
@@
The units of the GFDL data are $\text{kg m}^{-2}\text{s}^{-1}$, and for ERA5 m/h. For consistency, both are transformed to mm/d.
@@
  • only GFDL: rescaling the original 1°$\times$ 1.25° GFDL data to 1$\times$1° (64$\times$64 pixel).
  • Add +1 mm/d precipitation to each value in order to be able to apply a log-transformation to the data.
@@
  • Transform the data to the range [-1,1] in order to facilitate training convergence.
As part of the transformation $g$, the 1° GFDL data is bi-linearly upsampled. This and the downsampling and upsampling of ERA5 data, which is part of $f$, are already done during pre-processing. The downsampling of 0.25° ERA5 data (256$\times$256 pixel) to 1° (64$\times$64 pixel) is done by only keeping every fourth pixel in each field. For the just mentioned upsampling, we apply bi-linear interpolation to increase the resolution from 1° to 0.25°. Note that bi-linear interpolation to 0.25° does not increase the amount of information in the images compared to the 1° fields. After pre-processing the data as described, the embedding transformation $f$ is applied. The diffusion model is trained with the pre-processed $f(ERA5)$ as a condition and the original 0.25° ERA5 data as a target. Before we apply the embedding transformation $g$ we first pre-process the 1° GFDL data by applying quantile delta mapping (QDM [1]) with 500 quantiles. Bi-linear upsampling is afterward used to increase the resolution to 0.25$\times$0.25° (256$\times$256 pixels). The pre-processed data is used as input to the embedding transformation $g$. The corresponding output serves as the condition during the inference process of the diffusion model.
@@ 4.2 Embedding framework
Our framework introduces transformations $f$ $g$ that map OBS and ESM data to a shared embedding space $f:\mathbf{V^{obs}}\rightarrow\mathbf{V^{emb}}$ and $g:\mathbf{V^{esm}}\rightarrow\mathbf{V^{emb}}$. The goal is to do bias correction and downscaling of ESM fields, i.e., to obtain samples from the conditional distribution $\omega=p(OBS|ESM)$. Training a conditional model to approximate this distribution directly is not possible because $OBS$ and $ESM$ are unpaired. The arrows in the diagram of Figure 1 show that we can represent the mapping that achieves the bias correction and downscaling as $\omega=f^{-1}\circ g$. Our idea is to approximate $f^{-1}$ with a neural network $f^{-1}\approx\epsilon$. We chose a conditional diffusion model (DM), denoted by the conditional distribution $p(OBS|f(OBS))$, to approximate $f^{-1}=DM:\mathbf{V^{emb}}\rightarrow\mathbf{V^{obs}}$. The diffusion model (Fig. 1C) is only trained on pairs $(OBS,f(OBS))$. The shared embedding space allows us to evaluate the trained model on ESM embeddings $p(OBS|g(ESM))$, as all embeddings are identically distributed.
@@ 4.2.1 Constructing the embedding space
The goal of $f$ and $g$ is to map $OBS$ and $ESM$ to a shared embedding space, where $f(OBS)$ and $g(ESM)$ are identically distributed (Fig. 1). To achieve this, both embedded datasets need to be unbiased towards each other. $OBS$ and $ESM$ are biased towards each other in terms of statistical biases between the distributions and biases between small-scale patterns visible in the spatial power spectral density (PSD) (Fig S2A).
As mentioned earlier, the input to the embedding transformation $f$ is 0.25° ERA5 data, which is first downsampled, then upsampled and pre-processed. The input to the embedding transformation $g$ is the upsampled and pre-processed 0.25° GFDL data. By first downsampling ERA5 to 1° and then upsampling it to 0.25° we ensure that the fields match the information content of the original 1° GFDL fields.
To remove the small-scale pattern bias, we apply a noising procedure analogous to the forward diffusion process as part of $f$ and $g$. Both ERA5 and GFDL data are noised up to a certain spatial scale in the PSD, denoted by the point $s$. The scale $s$ is determined by the point where ERA5 and GFDL start to disagree in the PSD (Fig. 2), i.e., the intersection between the two. Adding noise ensures that $f(ERA5)$ is unbiased compared to $g(GFDL)$ in the PSD by erases all information beneath $s$. The transformations $f$ and $g$ utilize the same cosine scheduler as the forward diffusion process to add Gaussian noise to the data. ERA5 data undergoes 50 noise steps within $f$, while $g$ applies the same 50 noise steps to the GFDL data.
We ensure the observational and ESM data have aligned distributions by incorporating Quantile Mapping (QM) directly into the transformation $g$. It can directly be applied in pre-processing and only needs to be included in $g$. It’s important to clarify that QM is not included because the diffusion model is unable to do bias correction. QM is only used as a tool in our framework to ensure that in the embedding space $f(ERA5)$ and $g(GFDL)$ are identically distributed, such that $g(GFDL)$ can be used for the inference of the diffusion model.

### v2

@@ 4.1 Data
For the study region, we focus on the South American continent and the surrounding oceans. Specifically, the targeted area spans from latitude 0°N to 63°S and from longitude -90°W to -27°E. For the ablation study of the South Asian region, we selected an area from 0.75°N to 64.5°N latitude and from 42°E to 105.75°E longitude. The training period comprises ERA5 data from 1992-01-01 to 2011-01-01. The range of years included for the evaluation on ERA5 and GFDL spans from 2011-01-02 to 2014-12-01. Additionally, an extended 20-year window (1995–2014) is used for analyses requiring greater statistical robustness.
@@
The climate model output is taken from a state-of-the-art ESM from Phase 6 of the Coupled Model Intercomparison Project (CMIP6), namely GFDL-ESM4 [7]. We abbreviate the model with GFDL throughout the paper. The dataset contains daily precipitation data of the first ensemble member (r1i1p1f1) of the historical simulation (esm-hist). The data is available from 1850 to 2014, at 1° latitudinal and 1.25° longitudinal resolution and a daily temporal resolution.
GFDL-ESM4 [7] SSP5-8.5 represents a high-emission future pathway. We use daily-resolution data from the CMIP6 archive, provided at 1° latitude and 1.25° longitude spatial resolution, covering the period from 2015 to 2100.
MPI
For our ablation study, we repeat our experiments for the MPI-ESM HR model [12]. We abbreviate MPI-ESM-HR with MPI in the paper. The data has 0.9375°×0.9375° spatial resolution. We use daily data from 1992 to 2014 using data from 1992-2011 for training and 2011 to 2014 for inference.
@@
In order to benchmark our method, we first apply bilinear interpolation to increase the resolution of the GFDL fields from 1° to 0.25°. After that, we apply quantile delta mapping [3] to fit the upsampled GFDL data to the original 0.25° ERA5 data. QM is fitted on past observations and can then be used to correct the statistics of any (past/present) ESM field towards that reference period. We use quantile delta mapping (QDM) and chose the ERA5 training period from 1992-01-01 to 2011-01-01 as the reference period to fit the GFDL to ERA5. The benchmark dataset to evaluate our approach is then constructed by applying QM to the GFDL validation period (2011-01-02 to 2014-12-01). Some analyses required a longer evaluation period (1995-2014). To create a fair benchmark for these specific cases QDM was also recalibrated, it was both fitted and applied using data exclusively from this 1995-2014 window. For the SSP5-8.5 data, we use the 1995 to 2014 period of ERA5 as reference data and the historical GFDL data as the model input to fit the QDM. We then apply this mapping to the full time period of the GFDL SSP5-8.5 data (2015–2100).
@@
The units of the GFDL data and MPI data are $\text{kg m}^{-2}\text{s}^{-1}$, and for ERA5 $\text{m}\text{h}^{-1}$. For consistency, both are transformed to mm d-1.
@@
  • Only GFDL: rescaling the original $\times 1$.25° GFDL data to $\times1$° ($\times64$ pixel).
  • Only MPI: rescaling the original $\times 0.9375$° GFDL data to $\times1$° ($\times64$ pixel).
  • Add +1 mm d-1 precipitation to each value in order to be able to apply a log-transformation to the data.
@@
  • Transform the data to the range [-1,1] to facilitate the convergence of the training.
An ablation study (fig. S22) confirms the choice of our precipitation pre-processing pipeline, showing that omitting the log-transformation or the final range scaling leads to spectral discrepancies or distributional biases. As part of the transformation $g$, the 1° GFDL data is bilinearly upsampled. This and the downsampling and upsampling of ERA5 data, which is part of $f$, are already done during pre-processing. The downsampling of 0.25° ERA5 data ($\times256$ pixel) to 1° ($\times64$ pixel) is done by only keeping every fourth pixel in each field. For the just mentioned upsampling, we apply bilinear interpolation to increase the resolution from 1° to 0.25°. Note that bilinear interpolation to 0.25° does not increase the amount of information in the images compared to the 1° fields. After preprocessing the data as described, the embedding transformation $f$ is applied. The diffusion model is trained with the preprocessed $f(ERA5)$ as a condition and the original 0.25° ERA5 data as a target. Before we apply the embedding transformation $g$ we first pre-process the 1° GFDL data by applying quantile delta mapping (QDM [3]) with 500 quantiles. The bilinear upsampling is then used to increase the resolution to $\times0.25$° ($\times256$ pixels). The preprocessed data are used as input to the embedding transformation $g$. The corresponding output serves as the condition during the inference process of the diffusion model
@@ 4.2 Embedding framework
Our framework introduces transformations $f$ & $g$ that map OBS and ESM data to a shared embedding space $f:\mathbf{V^{obs}}\rightarrow\mathbf{V^{emb}}$ and $g:\mathbf{V^{esm}}\rightarrow\mathbf{V^{emb}}$. The goal is to do bias correction and downscaling of ESM fields, i.e., to obtain samples from the conditional distribution $\omega=p(OBS|ESM)$. Training a conditional model to approximate this distribution directly is not possible because OBS and ESM are unpaired. Therefore, we will train the model without the ESM data, only using OBS data and utilize a trick to enable transfer learning and inference on the ESM data. We apply transformations on ESM and OBS such that the resulting datasets are similarly distributed and therefore allow for generalization. The arrows in the diagram of Figure 1 show that we can represent the mapping that achieves the bias correction and downscaling as $\omega=f^{-1}\circ g$. Our idea is to approximate $f^{-1}$ with a neural network $f^{-1}\approx\epsilon$. We chose a conditional diffusion model (DM), denoted by the conditional distribution $p(OBS|f(OBS))$, to approximate $f^{-1}=DM:\mathbf{V^{emb}}\rightarrow\mathbf{V^{obs}}$. The diffusion model (Fig. 1C) is only trained on pairs $(OBS,f(OBS))$. The shared embedding space allows us to evaluate the trained model on ESM embeddings $p(OBS|g(ESM))$, as all embeddings are identically distributed.
@@ 4.2.1 Constructing the embedding space
The goal of $f$ and $g$ is to map OBS and ESM to a shared embedding space, where $f(OBS)$ and $g(ESM)$ are identically distributed (Fig. 1). To achieve this, both embedded datasets need to be unbiased towards each other. OBS and ESM are biased towards each other in terms of statistical biases between distributions and biases between small-scale patterns visible in the spatial power spectral density (PSD) (fig. S4A).
As mentioned earlier, the input for the embedding transformation $f$ is 0.25° ERA5 data, which is first preprocessed, then downsampled and upsampled. The input to the embedding transformation $g$ is the preprocessed and upsampled 0.25° GFDL data. By first downsampling ERA5 to 1° and then upsampling it to 0.25° we ensure that the fields match the information content of the original 1° GFDL fields.
To remove small-scale pattern bias, we apply a noising procedure analogous to the forward diffusion process as part of $f$ and $g$. Gaussian noise contains all frequencies in equal measure and the Fourier transform of Gaussian noise is itself Gaussian noise, so its power must be equal across all frequencies in expectation. The power spectrum of pure Gaussian noise corresponds to a horizontal line in the spectrum of Fig. 2A, reflecting the fact that it contains all frequencies in equal amounts. Adding noise to an image results in a hinge shape in the PSD of the noisy images (Fig. 2B, 2C and 2D). Increasing the variance of the noise increases its power and, as a result, its PSD will shift upward. Adding noise hence acts as a low-pass filter, while the variance of the added noise determines the cut-off frequency. Increasing variance leads to higher cut-off points as the power of the noisy frequencies increases. Both ERA5 and GFDL data are noised up to the cutoff frequency, denoted by $s$. The scale $s$ is determined by the point where ERA5 and the ESM data (in our case GFDL) start to disagree in their spatial PSDs (Fig. 2), i.e., the intersection between the two. Adding noise in this way ensures that $f(ERA5)$ is unbiased compared to $g(GFDL)$ in the PSD by erasing all information beneath $s$. In our implementation, the transformations $f$ and $g$ utilize the same cosine scheduler as the forward diffusion process to add Gaussian noise to the data. ERA5 data undergoes 50 noise steps within $f$, while $g$ applies the same 50 noise steps to the GFDL data. We ensure that the observational and ESM data have aligned distributions by incorporating Quantile Mapping (QM) directly into the transformation $g$. It only needs to be included in $g$. The quantile-mapped and bilinearly downscaled data is then noised as described above, as part of the embedding transformation. It is important to clarify that QM is not included because the diffusion model is unable to do bias correction. QM is only used as a tool in our framework to ensure that in the embedding space $f(ERA5)$ and $g(GFDL)$ are identically distributed, such that $g(GFDL)$ can be used for the inference of the diffusion model.